OpenStudy (anonymous):

what is the solution to the equation? sqrt2x+10 -6=2. Please show steps

OpenStudy (anonymous):

is the equation sqrt(2x + 10) - 6 = 2?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Can you help me out? Im so stuck

OpenStudy (anonymous):

since you have to solve for x, first add 6 to both sides so that you can isolate sqrt

OpenStudy (anonymous):

Okay sqrt(8x +10 -4=2? Im so bad at this

OpenStudy (anonymous):

Can you do a step by step @navk

OpenStudy (anonymous):

okay the equation is sqrt(2x + 10) - 6 = 2 or sqrt(8x +10) -4=2

OpenStudy (anonymous):

Okay, so how do we do this?

OpenStudy (anonymous):

@navk